Bob has twice as much money as Anthony.
Anthony has $7 more than Lisa.
Together they have $73.
How much does each have?

Answer by rapaljer(4671)

First, notice that Bob's money is given in terms of Anthony, and Anthony's money is given in terms of Lisa. So, let x = the quantity at the end of all of this, which is Lisa.

Let x = amount of money of Lisa
x+ 7 = amount of money of Anthony
2(x+7) = amount of money of Bob

The equation is that the sum of all these amounts = 73
x + x+7 + 2(x+7) = 73
x + x + 7 + 2x + 14 = 73
4x + 21 = 73
4x + 21 - 21 = 73 -21
4x = 52
x= $13 Lisa
x+ 7 = $20 Anthony
2(x+7) = $40 Bob

Check: 13+20 + 40 = 73
It checks!

R^2 at SCC